Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 02, 1996, pp. 29–33 Submitted: April 11, 1995. Published: September 9, 1996.

COMMON FIXED POINTS BY ISHIKAWA

ITERATES IN METRIC LINEAR SPACES

R.A.RASHWAN

Abstract. In this paper we obtain a common fixed point theorem for Ishikawa iterates of tw o self-mappings on a metric linear space under various contractive conditions.

(1991) A.M.S. Subject Classification Codes. 47H10.

Key Words and Phrases. Metric linear space, F - norm, Contractive mappings, Ishikawa iterates, Common fixed point.

Typeset by AMS-TEX

Mathematics Department, Faculty of Science, Assiut University, Assuit, EGYPT

E-mail Address: assiut@frcu.enu.eg

\begin{matrix}c \\ ◯ \end{matrix}1996 Cameron University